# L-MTP: Leap Multi-Token Prediction Beyond Adjacent Context for Large Language Models

### v1

@@ 6.2 LLM Inference Acceleration
There is a bunch of work focusing on accelerating LLMs, especially on their inference procedure [57, 58]. The remarkable techniques involve quantization [59, 60], pruning [61, 62], knowledge distillation [63, 44], compact architecture design [64, 65, 66], and dynamic network [67, 68]. The production deployment also advances to improve the inference efficiency, like memory management [69, 70] and parallelism [71, 72]. In this paper, we focus on the inference acceleration benefited by LLM decoding.
Prior works accelerate inference on greedy decoding [73, 74], while recent speculative decoding extends it with provably losslessness [37, 75, 76]. Speculative decoding follows the principle of draft-then-verify, where a draft model (smaller) efficiently generate multiple tokens for the parallel verification via the target model (larger). The drafting procedure can employ an independent model [38, 77] or enhance the own model [78], like adding additional FFN heads [73, 36, 40]. During the verification, the vanilla sampling only process tokens in a single draft sequence [37, 79], while recent methods utilize the token tree to verify multiple draft sequences in parallel [38, 39, 36, 40], further improve the token acceptance rate. In L-MTP, we apply our looking backward decoding by utilizing the additional LLM heads for self-speculative decoding, paired with the tree-based verification.
@@ Appendix A Detailed Proof of Theorem 3
[…]
@@
[…]
@@
$f(i+(i-1\text{ mod }k))\leq f(i),\text{with equality only when }i\text{ mod }k\equiv 1.$ (17)
@@
$\mathbf{\Delta}_{b}$ $=\sum_{m=1}^{n}\left[\prod_{i=1}^{m}\frac{1}{\exp[{\gamma(i+(i-1\text{ mod }k))}]}-\prod_{i=1}^{m}\frac{1}{\exp[{\gamma\cdot i}]}\right]+\sum_{m=n+1}^{kn-1}\prod_{i=1}^{m}\frac{1}{\exp[{\gamma(i+(i-1\text{ mod }k))}]}$ (21)
$=\sum_{m=1}^{n}\left[{\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod }k))}]}-{\exp[{-\gamma\sum_{i=1}^{m}i}]}\right]+\sum_{m=n+1}^{kn-1}{\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod }k))}]}.$ (22)
@@
$\mathbf{\Delta}_{b}$ $=\sum_{m=1}^{n}\left(\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod}2))]}-\exp[{-\gamma\sum_{i=1}^{m}i}]\right)+\sum_{m=n+1}^{2n-1}\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod }2))]}$ (23)
$=\sum_{m=1}^{n}\left(\exp[-\gamma\frac{m(m+1)}{2}+\left\lfloor\frac{m}{2}\right\rfloor]-\exp[-\gamma\frac{m(m+1)}{2}]\right)+\sum_{m=n+1}^{2n-1}\exp[{-\gamma(\frac{m(m+1)}{2}+\lfloor\frac{m}{2}\rfloor)}]$ (24)
@@
[…]
@@
[…]
@@
[…]
@@ B.1 Pseudo-Code for L-MTP
Training of L-MTP # multi-head forward computing $>$ for i in range(self.n_head): logits.append(self.heads[i](hidden_states)) # multi-head forward computing $<$ # .... # Leap Multi-token Prediction $>$ # if k == 1: L-MTP = MTP for i, logits_ in enumerate(logits): h_logits = logits_[:, : -(k*(i+1))].contiguous() h_labels = labels[..., k*(i+1) :].contiguous() loss_i = self.loss_fct(logits=h_logits, labels=h_labels, vocab_size=self.config.vocab_size) loss += loss_i # Leap Multi-token Prediction $<$ # ...
For inference, we add a cache to store the previous hidden state for our decoding. The decoding involves both previous and current hidden states to yield $k(n-1)$ additional tokens.
Inference of L-MTP # multi-head forward computing $>$ for i in range(self.n_head): logits.append(self.heads[i](hidden_states)) # multi-head forward computing $<$ # .... # get the previous one if self.model.past_hidden_states is None: self.model.past_hidden_states = hidden_states[:, -2].unsqueeze(1) past_hidden_states = torch.stack([self.model.past_hidden_states, hidden_states[:, -1].unsqueeze(1)], dim=0) logits_ = self.heads(past_hidden_states) lmtp_logits = logits_.flatten(start_dim=0, end_dim=1) # update the past hidden states self.model.past_hidden_states = past_hidden_states[-1] # ...
@@ B.4 Evaluation Benchmarks
MBPP & MBPP+ [47, 48] is a dataset of 974 Python programming problems designed to evaluate code generation and problem-solving abilities. Tasks range from simple functions to moderately complex algorithms, requiring correct implementation in Python. MBPP+ adds more unique test-cases (30$\times$) from the original MBPP [48].
HumanEval & HumanEval+ [49, 48] is a dataset of 164 hand-crafted Python programming problems, focusing on evaluating the functional correctness of code generation. Each problem includes a function signature, description, and test cases to verify the solution. HumanEval+ adds more unique test-cases (80$\times$) and fixes incorrect ground-truth solutions in HumanEval [48].
@@ Appendix D Additional Experimental Results
Training procedure. We present the training loss trends in Figure 11 for head warm-up and Figure 12 for full model tuning, with losses steadily decreasing and convergence to stability.

### v2

@@ 6.2 LLM Inference Acceleration
There is a bunch of work focusing on accelerating LLMs, especially on their inference procedure zhou2024survey ; linunlocking . The remarkable techniques involve quantization hubara2018quantized ; kim2023squeezellm , pruning ma2023llm ; gao2024disp , knowledge distillation hinton2015distilling ; peng2023instruction , compact architecture design child2019generating ; katharopoulos2020transformers ; hanhyperattention , and dynamic network rajbhandari2022deepspeed ; cao2025moe . The production deployment also advances to improve the inference efficiency, like memory management hooper2024kvquant ; liu2023scissorhands and parallelism oh2024exegpt ; mei2025helix . In this paper, we focus on the inference acceleration benefited by LLM decoding.
Prior works accelerate inference on greedy decoding stern2018blockwise ; sun2021instantaneous , while recent speculative decoding extends it with provably losslessness leviathan2023fast ; xia2024unlocking ; yin2024theoretical . Speculative decoding follows the principle of draft-then-verify, where a draft model (smaller) efficiently generate multiple tokens for the parallel verification via the target model (larger). The drafting procedure can employ an independent model miao2024specinfer ; yang2024multi or enhance the own model zhang2024accelerating , like adding additional FFN heads stern2018blockwise ; cai2024medusa ; li2024eagle . During the verification, the vanilla sampling only process tokens in a single draft sequence leviathan2023fast ; zhou2023distillspec , while recent methods utilize the token tree to verify multiple draft sequences in parallel miao2024specinfer ; sun2023spectr ; cai2024medusa ; li2024eagle , further improve the token acceptance rate. In L-MTP, we apply our looking backward decoding by utilizing the additional LLM heads for self-speculative decoding, paired with the tree-based verification.
@@ 7 Conclusion
Acknowledgements
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-002). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.
Xiaobo Xia is supported by MoE Key Laboratory of Brain-inspired Intelligent Perception and Cognition, University of Science and Technology of China (Grant No. 2421002). Xiu Su is supported by National Natural Science Foundation of China (No. 62406347). Shuo Yang is supported by the NSFC Young Scientists Fund (No. 62506096).
Appendix
@@ Appendix A Detailed Proof of Theorem 3
$\mathbb{E}[L]=\sum_{m=1}^{n}\Big(\prod_{i=1}^{m}\mathbb{E}_{\mathrm{x}_{\leq t}\sim\mathcal{D}}[p(\mathrm{x}_{t+i}|\mathrm{x}_{\leq t})]\Big)\quad\text{(Vanilla)}.$ (10)
@@
$\mathbb{E}[L]_{b}=\sum_{m=1}^{k(n-1)+1}\Big(\prod_{i=1}^{m}\mathbb{E}_{\mathrm{x}_{\leq t}\sim\mathcal{D}}[p(\mathrm{x}_{t+i}|\mathrm{x}_{\leq t{\color[rgb]{0,0.55,0.12}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.55,0.12}-(i-1)\text{ mod }k}})]\Big)\quad\text{(L-MTP)}.$ (11)
@@
[…]
@@
$\mathbf{\Delta}_{b}$ $=\sum_{m=1}^{n}\left[\prod_{i=1}^{m}\frac{1}{\exp[{\gamma(i+(i-1\text{ mod }k))}]}-\prod_{i=1}^{m}\frac{1}{\exp[{\gamma\cdot i}]}\right]$ (21)
$+\sum_{m=n+1}^{kn-1}\prod_{i=1}^{m}\frac{1}{\exp[{\gamma(i+(i-1\text{ mod }k))}]}$ (22)
$=\sum_{m=1}^{n}\left[{\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod }k))}]}-{\exp[{-\gamma\sum_{i=1}^{m}i}]}\right]$ (23)
$+\sum_{m=n+1}^{kn-1}{\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod }k))}]}.$ (24)
@@
$\mathbf{\Delta}_{b}$ $=\sum_{m=1}^{n}\left(\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod }2))]}-\exp[{-\gamma\sum_{i=1}^{m}i}]\right)$ (25)
$+\sum_{m=n+1}^{2n-1}\exp[{-\gamma\sum_{i=1}^{m}(i+(i-1\text{ mod }2))]}$ (26)
$=\sum_{m=1}^{n}\left(\exp[-\gamma\frac{m(m+1)}{2}+\left\lfloor\frac{m}{2}\right\rfloor]-\exp[-\gamma\frac{m(m+1)}{2}]\right)$ (27)
$+\sum_{m=n+1}^{2n-1}\exp[{-\gamma(\frac{m(m+1)}{2}+\lfloor\frac{m}{2}\rfloor)}]$ (28)
@@
$\leq\frac{\gamma}{2}\int_{0}^{n+1}x\exp[-\gamma\frac{x^{2}}{2}]dx$ (33)
$=\frac{\gamma}{2}\cdot\frac{1}{\gamma}\int_{0}^{\sqrt{\gamma}(n+1)}y\exp[-\frac{y^{2}}{2}]dy\quad(\text{let }y=\sqrt{\gamma}x)$ (34)
$=\frac{\gamma}{2}\cdot\frac{1}{\gamma}\left(1-\exp[-\gamma\frac{(n+1)^{2}}{2}]\right)\quad(\int_{0}^{a}ye^{-y^{2}/2}dy=1-e^{-a^{2}/2})$ (35)
@@
$\geq\int_{n+1}^{2n}\exp[-\gamma\frac{x^{2}}{2}]dx$ (39)
$\geq\frac{1}{\sqrt{\gamma}}\int_{\sqrt{\gamma}(n+1)}^{\sqrt{\gamma}\cdot 2n}\exp[-\frac{y^{2}}{2}]dy\quad(y=\sqrt{\gamma}x)$ (40)
@@
$\quad\quad(\int_{a}^{b}\exp[{-y^{2}/2}]dy\geq\frac{\exp[{-b^{2}/2}]}{b}-\frac{\exp[{-a^{2}/2}]}{a},\quad\text{for }b>a>0)$
@@ B.1 Pseudo-Code for L-MTP
Training of L-MTP # multi-head forward computing >>> for i in range(self.n_head): logits.append(self.heads[i](hidden_states)) # multi-head forward computing <<< # .... # Leap Multi-token Prediction >>> # if k == 1: L-MTP = MTP for i, logits_ in enumerate(logits): h_logits = logits_[:, : -(k*(i+1))].contiguous() h_labels = labels[..., k*(i+1) :].contiguous() loss_i = self.loss_fct(logits=h_logits, labels=h_labels, vocab_size=self.config.vocab_size) loss += loss_i # Leap Multi-token Prediction <<< # ...
For inference, we add a cache to store the previous hidden state for our decoding. The decoding involves both previous and current hidden states to yield k(n1)k(n-1) additional tokens.
Inference of L-MTP # multi-head forward computing >>> for i in range(self.n_head): logits.append(self.heads[i](hidden_states)) # multi-head forward computing <<< # .... # get the previous one if self.model.past_hidden_states is None: self.model.past_hidden_states = hidden_states[:, -2].unsqueeze(1) past_hidden_states = torch.stack([self.model.past_hidden_states, hidden_states[:, -1].unsqueeze(1)], dim=0) logits_ = self.heads(past_hidden_states) lmtp_logits = logits_.flatten(start_dim=0, end_dim=1) # update the past hidden states self.model.past_hidden_states = past_hidden_states[-1] # ...
@@ B.4 Evaluation Benchmarks
[…]
@@ Appendix D Additional Experimental Results
We present the training loss trends in Figure 11 for head warm-up and Figure 12 for full model tuning, with losses steadily decreasing and convergence to stability.
Table 3: The comparison to MTP with $n=4$ and $n=7$.
Math500 GSM8K MBPP MBPP+ HumanEval HumanEval+ MMLU IFEval Avg.
MTPn=4 25.40 45.79 67.72 57.67 65.85 59.15 65.21 35.49 52.79
MTP n=7 24.40 43.29 63.49 55.29 68.29 61.59 65.11 33.09 51.82
L-MTP k=2,n=4 28.20 46.25 67.99 59.26 67.68 60.37 65.23 35.01 53.75
L-MTP k=3,n=3 28.00 51.86 60.05 52.65 66.46 62.20 65.06 32.49 52.35
Comparison to MTP with $n=7$. Observed from Table 3, we can see that directly increasing the horizon of MTP does not improve the performance overall. But we can still observe some improvement on HumanEval. The interesting thing is that when we decrease the number of heads to 3, L-MTP ($k=3,n=3$) can achieve a better performance than MTP ($n=7$). The theory on the prediction analysis at different positions can answer this (Section 4). Distant tokens would lead to noise, while our leap can catch future tokens, but also leap some. An accumulated noise would be smaller than MTP. A smaller number of heads ($n=3$) while with the leaping strategy can achieve a comparable performance (52.79, MTP with $n=4$) or outperform MTP (51.82, $n=7$).